OpenStudy (anonymous):

IM confused about this question can someone explain it to me step by step please!!! -2(-c-12)=-2c-12

OpenStudy (anonymous):

First you need to distribute the -2. And don't for get that and neg x neg = pos.

OpenStudy (mathstudent55):

The first thing you need to do is use the distributive property on the left side. -2(-c - 12) = -2c - 12 -2(-c + -12) = -2c - 12 (-2)(-c) + (-2)(-12) = -2c - 12

OpenStudy (anonymous):

And then what im still confused

OpenStudy (anonymous):

What does that give you on the left side? and what does the equation look like now? show me please.

OpenStudy (anonymous):

who?

OpenStudy (anonymous):

Go ahead and type what you have, so far, in here.. then I can help you.

OpenStudy (anonymous):

Well the last thing i have now is (-2)(-c)+(-2)(-12)=-2c-12 which what mathstudent said but i had it long ago

OpenStudy (anonymous):

OK, so go ahead and simplify the left side.

OpenStudy (anonymous):

One sec imma do it over again cause im confused now

OpenStudy (anonymous):

and sorri the problem goes like -2(-c-12)=-2c-12

OpenStudy (anonymous):

Right, now distribute the -2 on the left side and then show me what the equation looks like at that point... and then I'll help you through the rest.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Omg ok im so confused over again can you show me what you mean because my head is spinning trying to figure out what you mean by -2 on the left side

OpenStudy (anonymous):

You multiply -2 x -c and then -2 x -12 .. that is how you do it. show me what you get please.

OpenStudy (anonymous):

Is this how you set it up -2 x -c-12=-2c-12 x -2?

OpenStudy (mathstudent55):

Multiply out the left side: (-2)(-c) + (-2)(-12) = -2c - 12 2c + 24 = -2c - 12

OpenStudy (anonymous):

The left side of mathstudent's reply is what you would get.

OpenStudy (mathstudent55):

Once you've multiplied out the left side, you need to bring all terms with c to the left side, and all numbers to the right side.

OpenStudy (anonymous):

OK ill try it

OpenStudy (mathstudent55):

2c + 24 = -2c - 12 In order to move -2c from the right side to the left side, you need to add 2c to both sides of the equation 2c + 24 = -2c - 12 +2c = +2c -------------------- (add both lines) 4c + 24 = - 12

OpenStudy (mathstudent55):

The next step is to move 24 to the right side. In order to do that, you need to subtract 24 from both sides. 4c + 24 = - 12 - 24 - 24 -------------- (add bnoth lines) 4c = -36

OpenStudy (anonymous):

Hmmm

OpenStudy (mathstudent55):

The last step is to divide both sides by 4 to get rid of the 4 that is multiplying c: 4c/4 = -36/4 c = -9 Final answer: c = -9

OpenStudy (anonymous):

Thanks that helped alot!!!

OpenStudy (mathstudent55):

You're welcome.